Help with these question. Please provide clear detailed explanation. Thanks!
media%2F100%2F1006e489-5417-4c13-992c-a6
Answer #1

Time period of Earth is 1 year. As Halley's comet crosses the Earth every 75 years, Its time period around the Sun is T=75\,yr

Time period T=2\pi\sqrt{\frac{R^3}{GM}}

Radius of orbit of Halley's comet is R=\left ( \frac{GM_ST^2}{4\pi^2} \right )^{1/3}

R=\left ( \frac{6.67*10^{-11}*(1.99*10^{30})(75*365*24*3600)^2}{4\pi^2} \right )^{1/3}

R=26.594*10^{11}\,m

Radius of comet's orbit is R=26.594*10^{11}\,m

b)

The closest the comet gets to Earth is R-R_E=26.594*10^{11}-1.5*10^{11}=25.094*10^{11}\,m

c)

Angular speed of comet is \omega=\frac{2\pi}{T}=\frac{2\pi}{(75*365*24*3600)}=2.657*10^{-9}\,rad/s

d)

Angular acceleration of comet is zero, since it is moving in a circle with uniform angular velocity.

c)